Decide whether these lines are parallel, perpindicular or neither
5x+2y=10
2x-5y=7
I think that are perpendicular because they are opposites!
Can you please help me?
Thank you

Answer by CharlesG2(834)

"Decide whether these lines are parallel, perpindicular or neither
5x+2y=10
2x-5y=7
I think that are perpendicular because they are opposites!
Can you please help me?
Thank you"
slope-intercept form of a line is y=mx+b,
where m is the slope,
and b is the y-intercept (vertical-intercept or the point (0,b))
5x + 2y = 10
2y = -5x + 10
y = (-5/2)x + 5 --> m = -5/2, b = 5
2x - 5y = 7
- 5y = -2x + 7
y = (2/5)x - (7/5) --> m = 2/5, b = -7/5
they are perpendicular
since the 2 slopes are reciprocals and their product is -1